OpenStudy (anonymous):

What is the product of the polynomials below? (2x2 + 5x + 5)(4x - 3)

OpenStudy (anonymous):

. 8x3 + 14x2 - 5x + 15 D. 8x3 + 14x2 + 5x - 15

OpenStudy (callisto):

(2x^2 + 5x + 5)(4x - 3) = 4x (2x^2 + 5x + 5) -3 (2x^2 + 5x + 5) =?

OpenStudy (marco26):

Use distributive property and simplify

OpenStudy (anonymous):

8x3 + 14x2 - 15x - 15

OpenStudy (callisto):

Nope ._.

OpenStudy (anonymous):

. 8x3 + 14x2 + 5x + 15

OpenStudy (callisto):

(2x^2 + 5x + 5)(4x - 3) = 4x (2x^2 + 5x + 5) -3 (2x^2 + 5x + 5) = 8x^3 + 20x^2 + 20x - 6x^2 - 15x -15 =?

OpenStudy (anonymous):

8x3 + 14x2 + 5x - 15

OpenStudy (callisto):

Yes :)